# Singing Voice Synthesis Using Differentiable LPC and Glottal-Flow-Inspired Wavetables

### v1

@@ 5.3 Training Configurations
We trained separate models for each singer, resulting in 8 models. The loss function is the summation of the multi-resolution STFT loss and f0 loss from SawSing with FFT sizes set to $\{512,1024,2048\}$, plus a binary cross entropy loss on voiced detection. We stopped the gradients from the harmonic source to the f0s and voice predictions to stabilise the training. We used Adam [41] for running all the optimisations. For DDSP and SawSing, the batch size and learning rate were set to 32 and 0.0005; for GOLF and PULF, the numbers were 64 and 0.0001. We used the ground truth f0s (extracted by WORLD [20]) for the harmonic oscillator of PULF during training due to stability issues. We trained all the models for 800k steps to reach sufficient convergence and picked the checkpoint with the lowest validation loss as the final model. [^3] The trained checkpoints, source codes, and audio samples are available at https://github.com/yoyololicon/golf.
@@ 6.1 Objective Evaluation
The objective metrics we choose are the multi-resolution STFT loss (MSSTFT), the mean absolute error (MAE) in f0, and the Fréchet audio distance (FAD) [42] on the predicted singing of the test set. Table 1 shows that DDSP has the lowest MSSTFT and f0 errors, while SawSing reaches the lowest FAD. GOLF and PULF show comparable results in f0 errors to other baselines. We report the memory usage when training these models and their real-time factor (RTF), both on GPU and CPU, in Table 2. The amount of memory required to train GOLF is around 35% of others, and it runs extremely fast, especially on the CPU.
@@ 7 Discussions
Given the evaluation results and the number of synthesis parameters in GOLF is roughly six times smaller than DDSP and SawSing, it is clear that GOLF’s synthesis parameters are a more compact representation. Comparing the differences between GOLF and PULF in Table 2, we can see that the performance gain is due to the use of wavetables. Other baselines synthesise band-limited harmonic sources with many sinusoids oscillating simultaneously, thus increasing the computational cost. PULF’s MOS score is much worse on f1, with noticeable artefacts in the unvoiced region and random components of the voice. After investigation, we found the noise gains $\beta_{n}$ predicted by PULF fluctuating at high speed, producing a modulation effect. This behaviour is also found in PULF trained on m1 and GOLF, even on the harmonic gains $\gamma_{n}$. Still, the amount of fluctuation is small and barely noticeable in the test samples. Given the available results, we could only conclude that this effect relates to the type of harmonic source and the range of f0, i.e., female singers have higher f0. This amplitude modulation effect cannot be observed on spectrograms and thus is not captured by the training loss we used. It could be an intrinsic drawback of using frame-wise LPC approximation, but more experiments and comparisons with sample-wise LPC are needed. In addition, SawSing got low scores for both singers because of the buzzing artefacts in the unvoiced region. This was pointed out in the original paper [18]. Although unvoiced gating (Sec. 3.4) reduces this problem to a large degree, human ears are susceptible to this effect. This could be an inherent problem in using a sawtooth as the harmonic source.
@@
The L2 loss shown in Table 2 demonstrates that GOLF matches phase-related characteristics more accurately than other models. Fig. 3 shows GOLF produces the most similar waveform to the ground truth. Other baselines’ waveforms are similar because they use the same additive synthesiser. It is possible to reduce their L2 loss by optimising the initial phases of the oscillators, but this cannot account for time-varying source shapes. Low L2 loss is a positive effect of the deterministic phase responses embedded in GOLF. This opens up many possibilities, such as decomposing and analysing the voice in a differentiable manner and training the vocoder using the time domain loss function. The latter could be a possible way to reduce the fluctuation problem discussed in the previous paragraph. The waveform matching of GOLF can be improved further by using a more flexible glottal source model, adding FIR and all-pass filters to account for the mixed-phase components of the voice.
@@
Lastly, we note that cascaded IIR filters provide an orderless representation (i.e. the cascading order does not affect the result). This results in the responsibility problem [44, 45] for the last layer of the encoder, which might be one of the reasons why GOLF and PULF are less stable to train than other baselines. Developing architectures that can handle orderless representation or switch to other robust representations are possible ways to address this.
@@ 8 Conclusions
We present a lightweight singing voice vocoder called GOLF, which uses wavetables with different glottal flows as entries to model the time-varying harmonic components and differentiable LPC filters for filtering both the harmonics and random elements. We show that GOLF requires less memory to train and runs an order of magnitude faster on the CPU than other DDSP-based vocoders but still attains competitive voice quality in subjective and objective evaluations. Furthermore, we empirically show that the predicted waveforms from GOLF represent the voice’s phase response more faithfully, which could allow us to use GOLF to decompose and analyse human voices.
@@ 9 Acknowledgements
The authors want to thank Moto Hira (mthrok) and Alban Desmaison (albanD) for reviewing our pull requests to Torchaudio. We are incredibly grateful to Parmeet Singh Bhatia (parmeet) for providing the first version of efficient IIR in the legendary lfilter.cpp file. We thank Ben Hayes for giving feedback on the equations of backpropagation through an IIR. The first author wants to exclusively thank Ikuyo Kita for giving positive, energetic support during the writing process. The first author is a research student at the UKRI Centre for Doctoral Training in Artificial Intelligence and Music, supported jointly by UK Research and Innovation [grant number EP/S022694/1] and Queen Mary University of London.

### v3

@@ 5.3 Training Configurations
We trained separate models for each singer, resulting in 8 models. The loss function is the summation of the multi-resolution STFT loss (MSSTFT) and f0 loss from SawSing with FFT sizes set to $\{512,1024,2048\}$, plus a binary cross entropy loss on voiced/unvoiced prediction. We stopped the gradients from the harmonic source to the f0s and voiced decisions to stabilise the training. We used Adam [41] for running all optimisations. For DDSP and SawSing, the batch size and learning rate were set to 32 and 0.0005; for GOLF and PULF, the numbers were 64 and 0.0001. We used the ground truth f0s (extracted by WORLD [20]) for the harmonic oscillator of PULF during training due to stability issues. We trained all the models for 800k steps to reach sufficient convergence and picked the checkpoint with the lowest validation loss as the final model [^3] The trained checkpoints, source codes, and audio samples are available at https://github.com/iamycy/golf..
@@ 6.1 Objective Evaluation
The objective metrics we choose are the MSSTFT, the mean absolute error (MAE) in f0, and the Fréchet audio distance (FAD) [42] on the predicted singing of the test set. Table 1 shows that DDSP has the lowest MSSTFT and f0 errors, while SawSing reaches the lowest FAD. GOLF and PULF show comparable results in f0 errors to other baselines. We report the memory usage when training these models and their real-time factor (RTF), both on GPU and CPU, in Table 2. The amount of memory required to train GOLF is around 35% of others, and it runs extremely fast, especially on the CPU.
@@ 7 Discussions
Given the evaluation results and the number of synthesis parameters in GOLF is roughly six times smaller than DDSP and SawSing, it is clear that GOLF’s synthesis parameters are a more compact representation. Comparing the differences between GOLF and PULF in Table 2, we can see that the performance gain is due to the use of wavetables. Other baselines synthesise band-limited harmonic sources with many sinusoids oscillating simultaneously, thus increasing the computational cost. PULF’s MOS score is much worse on f1, with noticeable artefacts in the unvoiced region and random components of the voice. After investigation, we found the noise gains $\beta_{n}$ predicted by PULF fluctuating at high speed, producing a modulation effect. This behaviour is also found in PULF trained on m1 and GOLF, even on the harmonic gains $\gamma_{n}$. Still, the amount of fluctuation is small and barely noticeable in the test samples. Given the available results, we could only conclude that this effect relates to the type of harmonic source and the range of f0, i.e., female singers have higher f0. This amplitude modulation effect cannot be observed in spectrograms and thus is not captured by the training loss we used. It could be an intrinsic drawback of using frame-wise LPC approximation, but more experiments and comparisons with sample-wise LPC are needed. In addition, SawSing produced low scores for both singers because of the buzzing artefacts in the unvoiced region. Although unvoiced gating (Sec. 3.4) reduces this problem to a large degree, human ears are susceptible to this effect. This could be an inherent problem in using a sawtooth as the harmonic source.
@@
The L2 loss shown in Table 2 demonstrates that GOLF matches phase-related characteristics more accurately than other models. Fig. 3 shows GOLF produces the most similar waveform to the ground truth. Other baselines’ waveforms are similar because they use the same additive synthesiser. It is possible to reduce their L2 loss by optimising the initial phases of the oscillators, but this cannot account for time-varying source shapes. Low L2 loss is a positive effect of the deterministic phase responses embedded in GOLF. This opens up many possibilities, such as decomposing and analysing the voice in a differentiable manner and training the vocoder using the time domain loss function. The latter could be a possible way to reduce the fluctuation problem discussed in the previous paragraph. The waveform matching of GOLF can be improved further by using a more flexible glottal source model, adding FIR and all-pass filters to account for the voice’s mixed-phase components and the recording environment’s acoustic response.
@@
Lastly, we note that cascaded IIR filters provide an orderless representation (i.e. the cascading order does not affect the outputs). This results in the responsibility problem [44, 45] for the last layer of the encoder, which might be one of the reasons why GOLF and PULF are less stable to train than other baselines. Developing architectures that can handle orderless representation or switch to other robust representations are possible ways to address this.
@@ 8 Conclusions
We present a lightweight singing voice vocoder called GOLF, which uses wavetables with different glottal flows as entries to model the time-varying harmonic components and differentiable LPC filters for filtering both the harmonics and random elements. We show that GOLF requires less memory to train and runs an order of magnitude faster on the CPU than other DDSP-based vocoders, but still attains competitive voice quality in subjective and objective evaluations. Furthermore, we empirically show that the predicted waveforms from GOLF represent the voice’s phase response more faithfully, which could allow us to use GOLF to decompose and analyse human voice.
@@ 9 Acknowledgements
The authors want to thank Moto Hira (mthrok), Christian Puhrsch (cpuhrsch), and Alban Desmaison (albanD) for reviewing our pull requests to TorchAudio. We are incredibly grateful to Parmeet Singh Bhatia (parmeet) for implementing the first version of efficient IIR in the TorchAudio codebase as lfilter.cpp. We thank Ben Hayes for giving feedback on the equations of backpropagation through an IIR. The first author wants to exclusively thank Ikuyo Kita for giving positive, energetic support during the writing process. The first author is a research student at the UKRI Centre for Doctoral Training in Artificial Intelligence and Music, supported jointly by UK Research and Innovation [grant number EP/S022694/1] and Queen Mary University of London.